OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -5 and a y-intercept of -1.

OpenStudy (ivettef365):

The equation of the line is y=mx+b where m = slope and b = y-int

OpenStudy (ivettef365):

can you take it from there

OpenStudy (anonymous):

Well actully i dont know, i kind of suck at math..

OpenStudy (anonymous):

And its not the math part, its the slopes i suck at.

OpenStudy (ivettef365):

you got that m= -5 and y-int = -1, then you have y = -5x -1 m= slope and b = y-int

OpenStudy (anonymous):

Hmm, So do i devide, add, multiply, or subtract?

OpenStudy (ivettef365):

The question is just asking for the equation of the line and the equation is y = -5x - 1 that is the answer

OpenStudy (anonymous):

Thanks, i owe you one!